### v1

# Response of wavelength-shifting and scintillating-wavelength-shifting fibers to ionizing radiation

Abstract
We report results of characterizing the response and light transport of wavelength-shifting (WLS) and scintillating-wavelength-shifting (Sci-WLS) fibers under irradiation by radioactive $\alpha$, $\beta$, and $\gamma$ sources. Light yield and light transmission were measured for the WLS fiber BCF-91A from Saint-Gobain and for a new Sci-WLS fiber EJ-160 from Eljen Technology.
The two variants with different fluor mixtures, EJ-160I and EJ-160II, exhibited approximately five and seven times higher light yield than BCF-91A, respectively, while their attenuation lengths were 3.80 m for BCF-91A, 4.00 m for EJ-160I, and 2.50 m for EJ-160II.
1 Introduction
Plastic wavelength-shifting and scintillating fibers offer efficient light collection and transport; thus, they have found a broad range of applications in particle and nuclear physics experiments. Recent examples include MINOS [1, 2], NOvA [3], T2K [4], the LHCb tracker [5], and the GERDA and LEGEND-200 liquid argon veto system [6, 7]. In the future, the LEGEND-1000 [8], the ePIC [9] and other experiments plan to continue using and refining such techniques.
In some detector designs, fibers are not only required to provide efficient light collection and wavelength-shifting, but also to meet additional stringent requirements on radiopurity [6, 7, 8]. An attractive option is the use of fibers as radiation detectors of natural nuclear and cosmogenic radioactivity. For example, fibers immersed in liquid argon can detect $\beta$ decays of 39Ar or $\beta$ decays of 42K from cosmogenic 42Ar [6, 7]. Additionally, properly doped and formulated fibers can self-tag their own radio-impurities. This may enhance the veto power while relaxing the radiopurity requirement of the manufacturing process.
We have partnered with Eljen Technology [10] to develop new fibers that would not only be competitive with the fibers available on the market, but would be better optimized for the needs of upcoming particle physics experiments such as LEGEND-1000 [8]. In this work, we report the results of a comprehensive study of two new scintillating-wavelength-shifting (Sci-WLS) fibers from Eljen Technology named EJ-160I and EJ-160II and compare their performance with the wavelength-shifting (WLS) fiber BCF-91A from Saint-Gobain [11], now Luxium Solutions [12].
2 Fiber samples
Table 1 and Figure 1 summarize the main physical and optical properties of three fibers: BCF-91A from Saint-Gobain and the newly developed EJ-160 fibers from Eljen Technology, which are produced in two variants—EJ-160I and EJ-160II—featuring different fluor mixtures. The BCF-91A fiber type was previously used in GERDA [6] and LEGEND-200 [7]. The test samples were supplied to us by a group at the Technical University of Munich. Figure 2 shows images of diamond fly-cut cross sections of fibers captured with a microscope [13]. The BCF-91A sample has a single cladding of approximately 0.03 mm thickness, while Eljen Technology fibers each have a cladding layer of 0.04 mm thickness.
Feature BCF-91A EJ-160
Variant standard EJ-160I EJ-160II
Type wavelength-shifting scintillating-wavelength-shifting
Cross-section 1 mm square 1 mm square
Cladding single(*) single(*)
Core material polystyrene polystyrene
Cladding material PMMA PMMA
Refractive index (core / cladding) 1.60/1.49 1.59/1.49
Cladding thickness 0.03 mm(*) 0.04 mm(*)
Table 1: Properties of tested fibers. PMMA is polymethyl methacrylate. (*) The thicknesses of claddings were measured by our microscope.
Figure 1: Absorption (left) and emission (right) spectra of the WLS fibers (manufacturers’ data). For reference, we include the emission spectra of tetraphenylbutadiene (TPB) from [14], which is often used for shifting scintillation light of liquid argon or liquid xenon, and the quantum efficiency of Silicon Photomultiplier (SiPM) S13360 from Hamamatsu Photonics [15]. All fibers and TPB spectra are normalized to their respective maxima.
Figure 2: Pictures of diamond fly-cut cross sections of the three tested fibers. These images were captured under a microscope with external illumination to highlight the core/cladding boundaries.
3 Experimental setup
The fibers were approximately 1.4 m long. They all had both ends optically coupled to SiPMs from Hamamatsu Photonics [15] using optical grease BC-630 from Saint-Gobain [11]. Figure 3 shows the SiPM mounted on a custom-designed board and presents its photon detection efficiency. SiPM outputs were digitized with a WaveRunner HRO 66Zi oscilloscope from Teledyne LeCory [16]. Representative traces and the corresponding ADC values histogram are shown in Figure 4.
We used standard $\alpha$, $\beta$, and $\gamma$ radioactive calibration sources to irradiate fibers at different distances from the end. SiPM pulse-heights were recorded on the scope and stored data were analyzed to evaluate the light-yield and light transmission of the fibers.
Figure 3: Left: SiPM readout board with the Hamamatsu Photonics SiPM S13360-3050CS coupled to a fiber. Fiber ends were polished using a diamond fly-cutter and coupled to the SiPM using optical grease. Right: Photon detection efficiency of the same SiPM [15].
Figure 4: Left: Typical SiPM signal traces recorded on the oscilloscope with a 0.5 photoelectron (p.e.) trigger threshold. Right: Histogram of SiPM ADC amplitudes obtained under LED excitation, yielding higher p.e. count. The legend shows the frequency of pulse generator and pulse width applied to the LED.
4 Fiber response measurements
4.1 Beta response
[…]
Figure 5: A schematic view of the setup for irradiation along a 138 cm-long fiber coupled to SiPM’s at both ends for $\beta$ and $\gamma$ irradiation studies.
Figure 6 presents the mean number of photoelectrons (p.e.) detected by SiPMs for the three fibers. A low statistical uncertainty of the mean is achieved by a 10-minute measurement at each source position, thus comprising tens of thousands of events. The left column of the figure, panels (a), (c) and (e), shows raw mean data, while the right column, panels (b), (d) and (f), shows the corresponding double-exponential fits. The fits follow the functional form:
$I=I_{\text{long}}\,e^{-x/\lambda_{\text{long}}}+I_{\text{short}}\,e^{-x/\lambda_{\text{short}}},$ (4.1)
where II is the number of photoelectrons (p.e.) recorded by the SiPM, and $x$ is the distance between the radiation source and the SiPM. $I_{\text{long}}$, $I_{\text{short}}$, $\lambda_{\text{long}}$, and $\lambda_{\text{short}}$ denote long and short components of the light yields and attenuation lengths. It is common in the literature and various applications that only $\lambda_{\text{long}}$ is used as “the fiber attenuation length”. However, at shorter distances a double-exponent function provides a much better description of the light yield [1, 2].
Figure 6: Light yields and fitted attenuation lengths of fibers irradiated with a 90Sr $\beta$ source.
For BCF-91A (Figure 6a), the blue, red, and green lines represent channel 1 (CH1) or SiPM left, channel 2 (CH2) or SiPM right, and their sum, respectively, exhibiting a symmetric dependence on the distance to the radioactive source. The light attenuation is well described by a double-exponential function.
By extrapolating the fits to the zero point of the x-axis (the source position relative to each SiPM), we estimate the light yield at $x=0$ for a given fiber, following a framework similar to that established in previous studies [17, 18].
They range from 12.7 p.e. for BCF-91A, 64.0 p.e. for EJ-160I, and 87.7 p.e. for EJ-160II. The long attenuation lengths that are fitted using formula 4.1 are 3.80 m (BCF-91A), 4.00 m (EJ-160I), and 2.50 m (EJ-160II). EJ-160II shows higher light yield at the cost of shorter attenuation lengths compared to the EJ-160I. In Figure 7 we directly compare the response of the three fibers to $\beta$ irradiation.
Figure 7: Summary of the $\beta$ response for the three fibers tested.
The systematic uncertainties of these measurements were initially evaluated by repeating the measurements “from scratch” (i.e., completely deconstructing and rebuilding the experimental setup) to project onto the overall systematic uncertainty of this technique. This included resetting the fiber on the SiPMs and the sources on the fibers tested. The statistical uncertainty is effectively negligible since each measurement contains tens of thousands of events.
4.2 Gamma response
[…]
Figure 8: (a) A schematic view of the setup for $\gamma$ testing. (b) A similar setup with narrower collimation. This setup was deemed impractical to collect large statistics. (c) Green and orange curves show measurement using a collimation scheme shown in a and b, respectively.
Figure 9: Light yield and fitted attenuation lengths of fibers irradiated with a 22Na $\gamma$ source.
The $\gamma$ irradiation results are shown in Figure 9. They mirror the $\beta$ trends. The left column of the figure, panels (a), (c) and (e), shows raw mean data, while the right column, panels (b), (d) and (f), shows the corresponding double-exponential fits. The light yields at the distance of 0 m from the SiPM, as extrapolated from the attenuation length fits, have values of 10.0 p.e. for BCF-91A, 55.8 p.e. for EJ-160I, and 76.9 p.e. for EJ-160II, respectively. In Figure 10 we directly compare the responses to $\gamma$ of the three fibers tested.
Figure 10: Summary of the $\gamma$ response for three tested fibers.
4.3 Alpha response
[…]
Figure 11: Schematic of $\alpha$ irradiation at multiple fiber lengths with single-end SiPM coupling.
As illustrated in Figure 12, the $\alpha$ irradiation results show similar trends as the $\beta$$\gamma$ irradiation. The panels (a), (b), and (c), shows raw mean data and its the corresponding double-exponential fits. The light yields extrapolated using fitted functions to the distance 0 m from the SiPM were 28.5 p.e., 81.6 p.e., and 102.9 p.e. for BCF-91A, EJ-160I, and EJ-160II, respectively. In Figure 13 we directly compare the $\alpha$ responses of the three fibers tested.
Figure 12: Light yield and fitted attenuation lengths of fibers irradiated with a 241Am $\alpha$ source.
Figure 13: Summary of the $\alpha$ response for three tested fibers.
4.4 Summary and discussion of measurements
Table 2 summarizes measurements discussed above. The table shows the number of photoelectrons extrapolated to position 0 m using fitted double-exponential functions for each fiber and irradiation type. These light yields are quoted with systematic uncertainties that were derived from uncertainties in the placement of radioactive sources and the positioning of fibers and their couplings to SiPMs. Statistical errors are negligible, as discussed previously.
For $\beta$ irradiation, the light yield of EJ-160I and EJ-160II is approximately 5.0 and 6.9 times higher than that of BCF-91A, respectively. For $\gamma$ irradiation, the corresponding factors are 5.6 and $7.$ For the $\alpha$ source, the enhancement is more modest, with EJ-160I and EJ-160II yielding factors of 2.9 and 3.6, respectively. This reduction may be attributed to the stronger quenching of the scintillation light for $\alpha$ particles in the polystyrene-based fibers EJ-160I and EJ-160II [20].
BCF-91A EJ-160I EJ-160II
Purpose WLS Scintillation and WLS Scintillation and WLS
Beta from 90Sr $(12.7\pm 0.8)$ p.e. $(64.0\pm 2.9)$ p.e. $(87.7\pm 3.7)$ p.e.
Gamma from 22Na $(10.0\pm 0.5)$ p.e. $(55.7\pm 2.4)$ p.e. $(76.9\pm 3.2)$ p.e.
Alpha from 241Am $(28.5\pm 2.2)$ p.e. $(81.6\pm 3.8)$ p.e. $(103\pm 11)$ p.e.
Table 2: Summary of light yield of fibers irradiated with $\alpha$, $\beta$, and $\gamma$ sources.
BCF-91A EJ-160I EJ-160II
Beta from 90Sr $\lambda_{\text{long}}$ = $(3.80\pm 1.25)$ m $\lambda_{\text{short}}$ = $(0.10\pm 0.02)$ m $\lambda_{\text{long}}$ = $(4.00\pm 2.62)$ m $\lambda_{\text{short}}$ = $(0.11\pm 0.01)$ m $\lambda_{\text{long}}$ = $(2.50\pm 1.49)$ m $\lambda_{\text{short}}$ = $(0.12\pm 0.01)$ m
Gamma from 22Na $\lambda_{\text{long}}$ = $(3.80\pm 0.82)$ m $\lambda_{\text{short}}$ = $(0.09\pm 0.01)$ m $\lambda_{\text{long}}$ = $(4.00\pm 2.39)$ m $\lambda_{\text{short}}$ = $(0.11\pm 0.01)$ m $\lambda_{\text{long}}$ = $(2.50\pm 1.31)$ m $\lambda_{\text{short}}$ = $(0.11\pm 0.01)$ m
Alpha from 241Am $\lambda_{\text{long}}$ = 3.80 m $\lambda_{\text{short}}$ = 0.07 m $\lambda_{\text{long}}$ = 4.00 m $\lambda_{\text{short}}$ = 0.09 m $\lambda_{\text{long}}$ = 2.50 m $\lambda_{\text{short}}$ = 0.09 m
Table 3: Summary of attenuation lengths of fibers irradiated with $\alpha$, $\beta$, and $\gamma$ sources. As discussed in the text, the values of $\lambda_{\text{long}}$ are taken from [21] due to the short lengths of tested fibers.
Table 3 summarizes the attenuation lengths obtained for each fiber and irradiation type. For long attenuation lengths, $\lambda_{\text{long}}$ is mainly determined by light collection data obtained at source distances of 50 cm or greater from the readout. At such distances, the number of detected photoelectrons is small, leading to relatively large uncertainties that propagate into the fitting of the extracted attenuation lengths. In a related study, but with 3 m long fibers illuminated by LEDs, a similar trend was observed [21]. Twice longer fibers resulted in much smaller uncertainties in attenuation lengths. To constrain the fit in the present analysis, we adopted $\lambda_{\text{long}}$ values determined from an LED-based study with longer fibers [21], thus we fit only $\lambda_{\text{short}}$. The resulting fits offered a reasonable description of the measured data, supporting the use of LED-based attenuation values as constraints in the fit.
We observe that although WLS fibers are not designed for scintillation, they produce detectable signals when irradiated by ionizing radiation. This is consistent with intrinsic scintillation in aromatic polymers such as polystyrene and polyvinyltoluene [22, 23].
5 Conclusions
We have characterized the radiation response of two new scintillating-wavelength-shifting fibers, EJ-160I and EJ-160II, under $\alpha$, $\beta$ and $\gamma$ irradiation, and compared their performance to that of the wavelength-shifting fiber BCF-91A. The EJ-160I and EJ-160II fibers yield approximately five and seven times more photoelectrons at the SiPM readout, respectively, compared to BCF-91A. The fitted attenuation lengths are 3.80 m for BCF-91A, 4.00 m for EJ-160I and 2.50 m for EJ-160II. When comparing the EJ-160 variants, EJ-160II shows higher light yield with a shorter attenuation length compared to the EJ-160I.
This publication presents partial results of the ongoing program to develop better Sci-WLS fibers and improve their radio purity. We also advance a comprehensive simulation framework to model light yield and photon transport in WLS and Sci-WLS fibers that will further benefit this work. A similar approach was previously validated for Kuraray Y-11 fibers [24]. Further fiber development and modeling results will be reported in future publications.
Acknowledgments
We thank Prof. S. Schönert, Dr. P. Krause, and the group at the Technical University of Munich for providing BCF-91A fiber samples and for insightful discussions. This work was supported in part by the University of Texas at Austin and the U.S. National Science Foundation under grant PHY-2312278.